Find the equivalent taxable yield of the municipal bond for tax brackets of zero, 10%, 20%, and 30%, if it offers a yield of 4.10%. (Round your answers to 2 decimal places.)

Tax Brackets
a. Zero %     
b. 10% %     
c. 20% %     
d. 30% %     
Answer #1

equivalent taxable yield=yield/(1-tax rate)

Tax brackets equivalent taxable yield
Zero 4.1/(1-0)=4.1%
10% 4.1/(1-0.1)=4.56%(Approx).
20% 4.1/(1-0.2)=5.13%(Approx).
30% 4.1/(1-0.3)=5.86%(Approx).